# On Mutation-Guided Unit Test Generation

### v2

@@ 4.3 Threats to Validity
Regarding the baseline approaches, we selected two representative methods: (1) EvoSuite, a widely-used state-of-the-art search-based testing tool, and (2)Genvanilla, a vanilla LLM-based prompt approach. These baselines allow us to evaluate whether incorporating mutation feedback in MutGen significantly improves fault detection capability, as measured by mutation score. Although we use Llama-3.3 as the base LLM in this study, our approach is not tied to any specific model. MutGen is model-agnostic by design, and evaluating its performance with other LLMs is left as future work.
@@
Randomness. Randomness may impact the effectiveness of EvoSuite, MutGen, and its variants. To mitigate this, and considering computational cost, we ran each tool on every subject three times and reported averaged results. We also performed statistical significance tests and calculated effect sizes, as detailed in Section 4.2.
@@
Data Leakage. While it is possible that some public code in benchmark datasets appears in the LLM’s training data (e.g., Llama-3.3), the specific mutants we generate are unlikely to be memorized. Moreover, we compare MutGen with a vanilla prompting baseline, and the observed performance gains suggest that our improvements stem from prompt engineering and strategy design rather than memorization.
5 Related Work
@@
Mutation-Guided Test Generation. Recent studies have questioned the sufficiency of coverage as an evaluation criterion. As shown in (Chekam et al., 2017; Andrews et al., 2006; Inozemtseva and Holmes, 2014), and pointed in recent work from Meta (Alshahwan et al., 2024; Foster et al., 2025), high coverage does not necessarily correlate with strong defect detection capability. Mutation testing offers a more rigorous alternative by assessing how well tests detect injected faults (mutants) (Chekam et al., 2017). Although some prior work (Dakhel et al., 2024; Foster et al., 2025) has attempted to improve mutation scores during test generation, it does not fully explore the underlying limitations of LLMs in detecting and killing live mutants. Our work aims to address this gap by incorporating mutation feedback into the prompt design process for LLMs, thereby enhancing the effectiveness of the generated test cases. In addition, we introduce an iterative generation process to push the limits of LLMs in killing mutants, providing insights into their current capabilities and boundaries. This exploration also highlights opportunities for future work on guiding LLMs to generate test cases with stronger defect detection capabilities.
@@
Other techniques. Automatic unit test generation has long been a central research topic in software engineering. Traditional efforts have largely focused on search-based and random testing techniques, with tools such as EvoSuite (Fraser and Arcuri, 2011) and Randoop (Pacheco and Ernst, 2007) significantly reducing the developers’ manual testing burden.
@@
With the advent of large language models (LLMs), a new paradigm has emerged in test generation. Numerous recent surveys have summarized the rapid advancements in this area (Fan et al., 2023; Wang et al., 2024; Hou et al., 2024); however, due to the large volume of related work, we only highlight a selection of representative examples in this section. Early research explored fine-tuning LLMs for code-related tasks. As LLMs became more powerful and general-purpose, however, prompt-based approaches gained traction, enabling test generation without task-specific training. Several studies (Siddiq et al., 2024; Yang et al., 2024; Yuan et al., 2024; Zhang et al., 2024; Li et al., 2023; Pizzorno and Berger, 2024) have investigated prompt engineering strategies and mechanisms to improve test case generation effectiveness using commercial LLMs, notably OpenAI’s GPT series (Achiam et al., 2023; Ouyang et al., 2022; Brown et al., 2020).
@@
Given the high cost of commercial LLMs, and the increasing availability of high-quality open-source alternatives such as the Llama and DeepSeek series (Grattafiori et al., 2024; Liu et al., 2024a), recent research has shifted toward optimizing prompt-based effectiveness for open-source models (Ouédraogo et al., 2024; Wang et al., 2025a; Deljouyi et al., 2024; Wang et al., 2025b; Huang et al., 2023; Kim et al., 2025). These models are becoming popular due to their accessibility and effectiveness in code-related tasks.
@@
LLM-based test generation has been applied across a wide range of testing scenarios. Some studies focus on multilingual test generation across languages such as Python (Wang et al., 2025a; Liu et al., 2024b; Xu et al., 2025; Tian et al., 2025), Java (Zhang et al., 2024; Wang et al., 2025a; Chen et al., 2024; Gu et al., 2024), and others (Wang et al., 2025a; Liu et al., 2024b; Zhang et al., 2025; Cheng et al., 2025; Wang et al., 2025c). Other works target specific testing goals, such as detecting particular bugs (Liu et al., 2024b; Foster et al., 2025), API testing (Cao et al., 2024; Deng et al., 2025; Li et al., 2025; Kim et al., 2025), or generating test oracles (Khandaker et al., 2025; Dinella et al., 2022; Binta Hossain and Dwyer, 2024). LLMs have also been integrated with traditional testing techniques, including mocking (Nan et al., 2025; ZHU et al., 2025), evolutionary algorithms (Taherkhani et al., 2024), and program analysis (Gu et al., 2025; Pan et al., 2025; Ryan et al., 2024; Pan et al., 2024). Evaluation in these works typically emphasizes code coverage metrics, such as branch and line coverage.
@@
Unlike prior work, our approach focuses on automatically improving the mutation score with LLMs, based on mutation testing feedback, since such score better reflects the fault-detection capability of test cases than code coverage (Chekam et al., 2017; Foster et al., 2025). As discussed in Section 2, high code coverage does not necessarily imply strong fault detection (Chekam et al., 2017). However, as demonstrated in our experiments (Section 4.2), improving mutation score often leads to increased code coverage as a by-product.
6 Discussion
@@
As described in Section 4 and 5, our experiments were conducted on independent, method-level subjects. This setup showed how well LLMs perform in killing mutants when not being confounded by broader program structures or dependencies. However, when transitioning to more complex datasets such as Defects4J, additional challenges arise, particularly in resolving dependencies to generate executable and meaningful test suites. We leave the investigation of such datasets for future work.
@@ 7 Conclusion
We evaluated MutGen on 104 subjects from the widely-used HumanEval-Java dataset and 100 subjects from Leetcode-Java, which we constructed from LeetCode algorithmic problem solutions. Experimental results show that MutGen significantly outperforms both EvoSuite and Llama-3.3 using a vanilla prompting strategy, achieving mutation scores of 89.5% and 89.1% on HumanEval-Java and Leetcode-Java, respectively.
@@
Acknowledgments
This work was supported by a research grant from Huawei Technologies and Research Ireland grant 13/RC/209. Lionel C. Briand’s contribution was partially funded by the Research Chair and Discovery Grant programs of the Natural Sciences and Engineering Research Council of Canada (NSERC).

### v4

@@ IV-C Threats to Validity
Regarding the baseline approaches, we selected two representative methods: (1) EvoSuite, a widely-used state-of-the-art search-based testing tool, and (2) Genvanilla, a vanilla LLM-based prompt approach. These baselines allow us to evaluate whether incorporating mutation feedback in MutGen significantly improves fault detection capability, as measured by mutation score. In addition to the above compared methods, we also compared MutGen with a variant of EvoSuite. Following published guidelines [31], we enabled the fitness function for strong mutation testing in EvoSuite. We conducted experiments on the two datasets using only the strong mutation fitness function. The average mutation scores for both datasets were 67.4% and 58.1%, respectively, which are similar to the results reported in our paper using the EvoSuite default setting. Although we use Llama-3.3 as the base LLM in this study, our approach is not tied to any specific model. MutGen is model-agnostic by design, and evaluating its performance with other LLMs is left as future work.
@@
Randomness. Randomness may impact the effectiveness of EvoSuite, MutGen, and its variant. To mitigate this, and considering computational cost, we ran each tool on every subject three times and reported average results. We also performed statistical significance tests and calculated effect sizes, as detailed in Section IV-B.
@@
Data Leakage. While some public code in benchmark datasets may appear in the LLM’s training data (e.g., Llama-3.3), the specific mutants we generate are unlikely to be memorized. Moreover, we compare MutGen with a vanilla prompting baseline, and the observed performance gains suggest that our improvements stem from prompt engineering and strategy design rather than memorization.
V Related Work
@@
Mutation-Guided Test Generation. Recent studies have questioned the sufficiency of coverage as an evaluation criterion. As shown in [18, 11, 39], and pointed in recent work from Meta [9, 28], high coverage does not necessarily correlate with strong defect detection capability. Mutation testing offers a more rigorous alternative by assessing how well tests detect injected faults (mutants) [18]. Although some prior work [23, 28] has attempted to improve mutation scores during test generation, it does not fully explore the underlying limitations of LLMs in detecting and killing live mutants. Our work aims to address this gap by incorporating mutation feedback into the prompt design process for LLMs, thereby enhancing the effectiveness of the generated test cases. In addition, we introduce an iterative generation process to push the limits of LLMs in killing mutants, providing insights into their current capabilities and boundaries. This exploration also highlights opportunities for future work on guiding LLMs to generate test cases with stronger defect detection capabilities.
@@
Other techniques. Automatic unit test generation has long been a central research topic in software engineering. Traditional efforts have primarily focused on search-based and random testing techniques, with tools such as EvoSuite [29] and Randoop [53] significantly reducing the developers’ manual testing burden. While EvoSuite was originally designed to maximize code coverage, prior work [31] has investigated enhancing mutation scores by modifying its fitness functions, which is an objective closely related to our work. We conducted experiments along these lines and discuss the results in Section IV.B. In short, results are not significantly different from those of the default setting of EvoSuite.
@@
With the advent of large language models (LLMs), a new paradigm has emerged in test generation. Numerous recent surveys have summarized the rapid advancements in this area [27, 66, 37]; however, due to the large volume of related work, we only highlight a selection of representative examples in this section. Early research explored fine-tuning LLMs for code-related tasks. As LLMs became more powerful and general-purpose, however, prompt-based approaches gained traction, enabling test generation without task-specific training. Several studies [61, 71, 72, 73, 45, 57] have investigated prompt engineering strategies and mechanisms to improve test case generation effectiveness using commercial LLMs, notably OpenAI’s GPT series [7, 52, 16].
@@
Given the high cost of commercial LLMs, and the increasing availability of high-quality open-source alternatives such as the Llama and DeepSeek series [33, 46], recent research has shifted toward optimizing prompt-based effectiveness for open-source models [51, 68, 24, 69, 38, 43]. These models are becoming popular due to their accessibility and effectiveness in code-related tasks.
@@
LLM-based test generation has been applied across a wide range of testing scenarios. Some studies focus on multilingual test generation across languages such as Python [68, 47, 70, 63], Java [73, 68, 20, 35], and others [68, 47, 74, 21, 67]. Other works target specific testing goals, such as detecting particular bugs [47, 28], API testing [17, 25, 44, 43], or generating test oracles [42, 26, 15]. LLMs have also been integrated with traditional testing techniques, including mocking [49, 76], evolutionary algorithms [62], and program analysis [34, 56, 60, 55]. Evaluation in these works typically emphasizes code coverage metrics, such as branch and line coverage.
@@
Unlike prior work, our approach focuses on automatically improving the mutation score with LLMs, based on mutation testing feedback, since such a score better reflects the fault-detection capability of test cases than code coverage [18, 28]. As discussed in Section II, high code coverage does not necessarily imply strong fault detection [18]. However, as demonstrated in our experiments (Section IV-B), improving mutation score often leads to increased code coverage as a by-product.
VI Discussion
@@
As described in Section IV and V, our experiments were conducted on independent, method-level subjects. This setup showed how well LLMs perform in killing mutants when not being confounded by broader program structures or dependencies. However, when transitioning to more complex datasets such as Defects4J, additional challenges arise, particularly in resolving dependencies to generate executable and meaningful test suites. We leave the investigation of such datasets for future work.
@@ VII Conclusion
We evaluated MutGen on 104 subjects from the widely-used HumanEval-Java dataset and 100 subjects from Leetcode-Java, which we constructed from LeetCode algorithmic problem solutions. Experimental results show that MutGen significantly outperforms both EvoSuite and Llama-3.3 using a vanilla prompting strategy, achieving mutation scores of 89.5% and 89.1% on HumanEval-Java and Leetcode-Java, respectively. Since MutGen aims to maximize the mutation score of generated test cases, we also compare it with a variant of EvoSuite (EvoSuitemut), where the fitness function of maximizing coverage is replaced with one that maximizes mutation score. The results show that EvoSuitemut performs similarly to EvoSuite, but both are outperformed by MutGen.